What is 24 to the third power
1
Expert's answer
2019-11-15T12:04:56-0500

243=(23*3)3=(23)3*33=23*3*33=29*33={22=2*2=4, 23=2*2*2=4*2=8, 24=2*2*2*2=8*2=16, 25=2*2*2*2*2=16*2=32, 26=2*2*2*2*2*2=32*2=64, 27=2*2*2*2*2*2*2=64*2=128, 28=2*2*2*2*2*2*2*2=128*2=256, 29=2*2*2*2*2*2*2*2*2=256*2=512, 32=3*3, 33=3*3*3=9*3=27}=512*27=512*(20+7)=512*20+512*7=512*2*10+512*7=(500+12)*2*10+(500+12)*7=(500*2+12*2)*10+500*7+12*7=(1000+24)*10+3500+84=1024*10+3584=10240+3584=13824